OpenStudy (anonymous):

Gabrielle is 9 years younger than Mikhail. The sum of their ages is 81 . What is Mikhail's age?

OpenStudy (anonymous):

Let g = gabrielle's age Let m = Mikhail's age. You can make two equations and then solve them using substitution: Gabrielle is 9 years younger than Mikhail, so: g = m - 9 The sum of their ages is 81, so: m + g = 81 Substitute the first equation into the second equation: m + (m - 9) = 81 m + m - 9 = 81 2m - 9 = 81 2m = 90 Can you solve for m?